# and Tight DP Auditing

### v2

@@ Abstract
This paper systematizes research on auditing Differential Privacy (DP) techniques, aiming to identify key insights into the current state of the art and open challenges. First, we introduce a comprehensive framework for reviewing work in the field and establish three cross-contextual desiderata that DP audits should target—namely, efficiency, end-to-end-ness, and tightness. Then, we systematize the modes of operation of state-of-the-art DP auditing techniques, including threat models, attacks, and evaluation functions. This allows us to highlight key details overlooked by prior work, analyze the limiting factors to achieving the three desiderata, and identify open research problems. Overall, our work provides a reusable and systematic methodology geared to assess progress in the field and identify friction points and future directions for our community to focus on.
@@ 1 Introduction
In today’s data-driven economy, a large number of potentially sensitive datasets are collected, shared, and released, including demographic [107], behavioral [106], and medical [86] data. Alas, this comes with inherent privacy risks that are difficult to mitigate in practice [103, 82]. In this context, Differential Privacy (DP) [38] provides a robust mathematical framework offering formal privacy guarantees. More precisely, DP bounds the impact that any single user can have on the output of a data release. In recent years, DP techniques have been deployed in both the public and private sectors. High-profile examples include the 2020 US Census release [2], Apple’s local DP system for learning user preferences [9], Google’s next-word prediction models [75], and Microsoft’s synthetic data release with the UN’s International Organization for Migration [78].(For a comprehensive and up-to-date list of real-world DP deployments, we refer the reader to [31].)
@@
DP auditing is also a powerful and versatile technique to investigate the tightness of theoretical guarantees [53, 84, 83, 7, 72], develop theoretical intuitions for improved privacy analyses [85, 19], and refute conjectures [5]. Moreover, it can be used to bound the adversarial success with respect to privacy attacks like membership inference [53], attribute inference [73], or reconstruction [50, 72].
Motivation. To the best of our knowledge, research on DP auditing dates back to 2013 [34]; yet, extending and improving DP auditing techniques or proposing new ones remains an active area of research today. Early work [33, 18] required running the implementations millions of times, thereby making their computational overhead prohibitively high for auditing complex private machine learning pipelines. More recent efforts have focused on efficiency, aiming to reduce the number of implementation runs down to a single one [101, 72, 112]. However, for complex DP mechanisms, these techniques are still largely ineffective in identifying bugs and violations as the audits fail to produce estimates of the empirical privacy leakage that are “close” to the theoretical DP bounds—i.e., the audits are not “tight.”
@@
Roadmap. In this paper, we set out to systematize the large body of recent work on DP auditing along several key axes, aiming to identify fresh insights, open research problems, and promising directions to guide future work in improving auditing techniques. We define and formalize three desiderata for DP auditing – efficient, end-to-end, and tight – and provide a comprehensive framework for evaluating progress in the field. Finally, we set out to systematize the modes of operation of current DP auditing techniques, including threat models, attacks, and evaluation functions.
Our analysis sheds light on key details overlooked by prior work and the limiting factors in achieving the three desiderata. For instance, one of the main limitations to tighter audits is the lack of tight theoretical privacy analyses for sub-sampling schemes, such as shuffling. We also show that, while recent audits have focused on the restrictive “black-box” threat model (where adversaries can only insert a single input sample and observe only the algorithm’s final output), much less attention has been given to slightly stronger threat models where audits are equally not tight. Finally, we find that, while recent work achieves efficient and tight audits for simple mechanisms, this remains far from reach, even in strong threat models, for more realistic and complex mechanisms like DP-SGD [1].
Contributions. In summary, our work makes three key contributions:
@@
We analyze the progress made by work in this field vis-à-vis three main desired properties, i.e., efficiency, end-to-end-ness, and tightness, underscoring the importance of achieving all three simultaneously both for simple mechanisms and complex private machine learning pipelines.
@@ 2 Background
In this section, we review background concepts related to Differential Privacy. Readers familiar with these topics can skip it without losing continuity.
@@ 2.4 Hypothesis Testing Interpretation of DP
The notion of hypothesis testing interpretation of DP, introduced in [57] for ()-DP, allows to evaluate a mechanism’s privacy guarantees based on the difficulty an adversary faces in distinguishing between two adjacent datasets. Given a random output from a mechanism (where for adjacent ), DP can be interpreted as inducing the following a hypothesis test:
Privacy Region. For a choice of rejection region , i.e., is rejected if , the type I and type II errors are defined as and , respectively. Thus, ()-DP bounds the type I and type II errors as per the following theorem:
@@ 3 DP Auditing
We now discuss the notion of DP auditing – a set of empirical techniques involving an adversary instantiated against a DP mechanism to distinguish between the mechanism’s outputs on adjacent datasets. We discuss its two main motivating use cases and the desired properties we believe DP auditing algorithms should satisfy.
@@ 3.1 Goals of DP Auditing
3.1.1 Verifying Privacy Guarantees
While the increased adoption of DP in real-world applications [2, 9, 78, 75, 63, 88] is a promising development for privacy-enhancing technologies, it can often be difficult to implement DP correctly. Bugs leading to unintended privacy leakage, in some cases completely breaking DP guarantees, have been found in many implementations [70, 118, 55, 67, 8, 6, 46]. For instance, Lyu et al. [70] and Zhang et al. [118] focus on the differentially private Sparse Vector Technique (SVT) [39] and show, through manual analysis, that most of its implementations [96, 40, 65, 102, 22, 70] are not differentially private at all.
@@ 3.1.1 Verifying Privacy Guarantees
However, this approach faces some fundamental limitations. First, type systems implicitly trust that some “basic” DP algorithms (e.g., Gaussian mechanism, Laplace mechanism) are implemented correctly by the underlying language. As pointed out in [67], this may not necessarily be the case if, e.g., the noise addition mechanisms suffer from floating-point vulnerabilities [79]. Second, these type systems require mechanisms to be implemented in a very specific language, thus, they are substantially limited to simple algorithms like SVT or Report Noisy Max [40]. Specifically, they struggle to capture the complexities involved in algorithms like RAPPOR [42] or DP-SGD [1].
Enter DP Auditing. Rather than statically checking program specifications, DP auditing allows an auditor to run experiments and statistically refute the theoretical privacy guarantees provided by DP. In other words, if the auditor fails, programmers and clients can be assured that the implementation satisfies the DP guarantees at some (possibly pre-defined) level of confidence—e.g., for more critical applications, the auditor can perform the test at 99.999% confidence [105] or 95% for less critical applications [53].
Overall, DP audits can offer more flexibility than static methods, accommodating more complex algorithms such as private machine learning pipelines [53]. They are also easier to implement as they do not involve specific programming languages or typing rules.
3.1.2 Empirical Privacy Estimation
DP auditing can also be useful to provide precise estimates of the privacy leakage in different settings, e.g., vis-à-vis different real-world adversaries and/or attacks.
@@ 3.2 DP Auditing Algorithms
This definition also encodes a notion of “correctness” for DP auditing algorithms, which informally states that the output of the algorithm agrees with the actual guarantees satisfied by the mechanism (i.e., if satisfies -DP, the algorithm confirms this to be true as well). However, this definition admits vacuous algorithms (e.g., an algorithm that always outputs ‘1’); thus, in Section 3.3, we define additional desirable properties of DP auditing algorithms.
@@
First, the Adversary chooses an initial dataset in the data domain and distributions from which special data points used to audit the mechanism, i.e., canaries, are sampled from. The Adversary can also provide additional “auxiliary” information to the mechanism if necessary. Next, the Challenger runs the mechanism times, each time sampling fresh canaries from the canary distributions to form the input dataset. The Challenger then sends the outputs back to the Adversary. The Adversary post-processes the outputs of the mechanism, typically using a privacy attack . This is because the outputs of the mechanism may be in an arbitrary domain , which can be difficult to design a decision function around (e.g., to decide if the mechanism satisfies the claimed guarantees). Thus, the post-processing yields simple observations that can be easily analyzed.
Finally, the audit’s outcome can be computed from the observations using an evaluation function , which outputs either a bit for the verification of a theoretical DP guarantee or for empirical privacy estimation. This is typically done by relying on the hypothesis testing interpretation of DP discussed above. In practice, the outcome of the audit is typically computed with a desired level of confidence , especially when refuting/confirming the claimed DP guarantees of the mechanism.
@@ 3.3 Desiderata: The Gold Standard of DP Auditing
Next, we define three key properties we believe audits should satisfy – efficiency, end-to-end-ness, tightness. We believe these properties are crucial for the widespread and rigorous adoption of DP auditing, especially in production systems. As discussed later, achieving all three simultaneously mostly remains an open research problem.
@@
A DP auditing algorithm is efficient if it requires the mechanism to be run only a few times (typically once or twice).
End-to-End. DP audits should also apply to end-to-end privacy guarantees of the implementations; ideally, these should be run by simply providing inputs to the mechanism and observing its final output. This is often referred to as “black-box” auditing [66] and reflects realistic threat models encountered in practical settings as opposed to worst-case threat models assumed by the privacy analysis of certain mechanisms. For instance, the privacy analysis of the DP-SGD [1] algorithm assumes that all intermediate models are released, even though in practice only the final model may be accessible to the adversary. Since “black box” may have different meanings in different areas (e.g., query-access to machine learning models [100] vs. access to fixed generated synthetic data records [51], etc.), we use the notion of “end-to-end” auditing to avoid confusion.
End-to-end audits are also broadly applicable to many mechanisms as they do not require access to how they are implemented internally. For example, although different DP mechanisms can be used to train private ML models (e.g., DP-SGD [1], PATE [90], DP-ZO [104]), relying on one single auditing technique that takes in input the training dataset and observe the trained model’s outputs would make it significantly easier to integrate auditing into different training pipelines. Once again, this would also more closely match the capabilities of realistic adversaries.
@@
Tightness. Informally, a tight audit means that the empirical audit is consistent with the theoretical privacy analysis. More precisely, the empirical estimates obtained through auditing should not be far from the theoretical guarantees (e.g., and ).
Moreover, when verifying DP guarantees, tight DP audits would ensure that implementations are successfully verified only if their actual privacy leakage is strictly lower than the theoretical DP bounds. As a consequence, tight audits are needed in this context to reinforce their accuracy and provide strong confidence in the process. By contrast, a loose audit that successfully verifies implementations where the actual privacy leakage is, say, twice the theoretical upper bounds would inherently not be very useful to identify bugs.
@@
4 The DP Auditing Body of Work
In this section, we present the methodology we use to compile a comprehensive and representative snapshot of research on DP auditing.
@@ 4 The DP Auditing Body of Work
Final List. After manual review, we identified 45 relevant DP auditing papers, which form the basis for our systematization effort. To ease presentation, in Table 1, we report 21 key papers (out of the 45) that arguably represent the state of the art, while reporting earlier/seminal work in Appendix B. Specifically, Table 1 lists papers that are the first to audit a particular DP guarantee (e.g. [33, 83]), mechanism (e.g. [53, 73]), or achieve the current “best” auditing results (e.g., [83, 112]). Along with the relevant selected papers, we also report key details of the auditing methods introduced in these papers, which we extract through our systematization process as discussed later in Section 5.
Table 1 reveals that state-of-the-art work in DP auditing spans a wide range of mechanisms audited, from “simple” (e.g., Laplace, Noisy Histogram, etc.) to more complex machine learning and federated learning pipelines, vis-à-vis various DP guarantees and adjacency notions and employing different sub-sampling techniques. Thus, to support this wide range of mechanisms and techniques, auditing algorithms use different privacy attacks and evaluation functions, which we analyze in Sections 5.2 and 5.3, respectively. Overall, DP auditing techniques have achieved the three desiderata discussed in Section 3.3 only for simple mechanisms (e.g., Gaussian), while doing so for complex mechanisms remains an elusive task.
Although our selection was done in Jan 2025, we are confident it is meaningfully representative of the state of the art in DP auditing as of the time of submission; i.e., even if one were to add or remove a small number of papers to our list, our main takeaways would not substantially change.
Out-of-scope Research. As our systematization focuses on DP auditing, we did not include work estimating privacy leakage from algorithms using privacy metrics other than DP guarantees, e.g., using attack success rate and/or adversarial advantage [93, 54]. We also excluded research focusing on using static methods, such as type checkers, to verify the DP guarantees of an implementation without executing the program. While these methods are useful in various ways, they are arguably limited in the range of mechanisms they can audit, which makes them ineffective for verifying the DP guarantees of complex modern DP mechanisms, e.g., in the context of deep learning, as discussed in Section 3.1.

### v3

@@ Abstract
In this paper, we systematize research on auditing Differential Privacy (DP) techniques, aiming to identify key insights and open challenges. First, we introduce a comprehensive framework for reviewing work in the field and establish three cross-contextual desiderata that DP audits should target—namely, efficiency, end-to-end-ness, and tightness. Then, we systematize the modes of operation of state-of-the-art DP auditing techniques, including threat models, attacks, and evaluation functions. This allows us to highlight key details overlooked by prior work, analyze the limiting factors to achieving the three desiderata, and identify open research problems. Overall, our work provides a reusable and systematic methodology geared to assess progress in the field and identify friction points and future directions for our community to focus on.
@@ 1 Introduction
Aiming to mitigate the privacy risks stemming from sharing or releasing sensitive datasets [107, 86], Differential Privacy (DP) [40] provides a robust mathematical framework offeringăformal privacy guarantees. More precisely, DP bounds the impact that any single user can have on the output of a data release. In recent years, DP techniques have been deployed in both the public and private sectors. High-profile examples include the 2020 US Census release [2], Apple’s local DP system for learning user preferences [10], Google’s next-word prediction models [78], and Microsoft’s synthetic data release with the UN’s International Organization for Migration [82].111Refer to https://desfontain.es/blog/real-world-differential-privacy.html for a comprehensive list of real-world DP deployments.
@@
DP auditing is a powerful and versatile technique to investigate the tightness of theoretical guarantees [55, 88, 87, 8, 75], develop theoretical intuitions for improved privacy analyses [89, 21], and refute conjectures [6]. Moreover, it can be used to bound the adversarial success with respect to privacy attacks like membership inference [55], attribute inference [76], or reconstruction [52, 75]. Overall, DP auditing has important connections to many areas of privacy and security research, including DP verification, privacy accounting, etc.
Motivation. While research on DP auditing dates back to at least 2013 [35], extending and improving DP auditing techniques or proposing new ones still is a very active area of research. Early work [34, 19] required running the implementations millions of times, making their computational cost prohibitively high for auditing complex private machine learning pipelines. Recent efforts have focused on efficiency, aiming to reduce the number of implementation runs down to a single one [105, 75, 116]. However, for complex DP mechanisms, these techniques are largely ineffective in identifying bugs and violations as the audits fail to produce estimates of the empirical privacy leakage that are “close” to the theoretical DP bounds—i.e., the audits are not “tight.”
@@
Roadmap. In this paper, we undertake a systematic examination of the recent (and growing) literature on DP auditing across several key dimensions. Our objective is to distill new insights, highlight open research challenges, and delineate promising directions to improve auditing methodologies. We formalize three desiderata for DP auditing: efficiency, end-to-end-ness, and tightness, and present a comprehensive framework based on them to assess progress within the field. Finally, we provide a structured systematization of the operational modes of existing DP auditing techniques, encompassing threat models, attack strategies, and evaluation functions.
Our analysis highlights important aspects that have been overlooked by prior work, as well as the key obstacles to achieving the three desiderata. For instance, a major limitation to tighter audits is the absence of tight theoretical privacy analyses for sub-sampling schemes, such as shuffling. We also show that, although recent audits have largely concentrated on the restrictive “black-box” threat model (where adversaries can only insert a single input sample and observe only the algorithm’s final output), far less attention has been devoted to slightly stronger threat models, in which audits also fail to be tight. Finally, while efficient and tight audits have recently been presented for simple mechanisms, extending these guarantees to more realistic and complex mechanisms like DP-SGD [1] remains a distant goal even under stronger threat models.
Summary of Contributions. In short, in this work:
1.
@@
2.
We analyze the progress made by work in this field vis-à-vis three main desired properties, i.e., efficiency, end-to-end-ness, and tightness, which are important considerations for data and privacy practitioners when choosing DP auditing techniques to deploy.
3.
@@ 2.4 Hypothesis Testing Interpretation of DP
The notion of hypothesis testing interpretation of DP, introduced in [59] for ()-DP, allows to evaluate a mechanism’s privacy guarantees based on the difficulty an adversary faces in distinguishing between two adjacent datasets. Given a random output from a mechanism (where for adjacent ), DP can be interpreted as inducing the following hypothesis test:
Privacy Region. For a choice of rejection region , i.e., is rejected if , the type I (False Alarm) and type II (Missed Detection) errors are defined as and , respectively. Thus, ()-DP bounds the type I and type II errors as per the following theorem:
@@
2.5 Related Work
To the best of our knowledge, our work is the first to systematically review and analyze the body of work on DP auditing. In the following, we briefly review other systematization-of-knowledge and survey papers on related topics. Namatevs et al. [85] recently present a survey of DP auditing research; however, they do not analyze relevant work within a systematic framework. By contrast, we do so to comprehensively analyze the state of the art, measure progress, identify gaps and open research problems primarily along three key characteristics: whether audits are efficient, end-to-end, and tight.
Systematization of knowledge efforts in related areas includes the work by Papernot et al. [95], who present attacks and defenses on (non-private) machine learning systems using a comprehensive adversarial framework. Salem et al. [101] focus on privacy risks and provide a unifying framework relying on privacy games. Finally, Desfontaines and Pejó [33] systematize the various variants and extensions of DP. While DP auditing relies on and is informed by the notions of privacy attacks, games, and DP variants, our work focuses on the specific process of deriving empirical bounds, which has thus far not been explored in prior work.
@@ 3 DP Auditing
DP auditing denotes a set of empirical techniques involving an adversary instantiated against a DP mechanism to distinguish between outputs on adjacent datasets. We now discuss its two main motivating use cases and the desired properties we believe DP auditing algorithms should satisfy.
@@ 3.1 Goals of DP Auditing
Verifying Privacy Guarantees. While increased adoption of DP in real-world applications [2, 10, 82, 78, 66, 92] is a promising development for privacy-enhancing technologies, it can often be difficult to implement DP correctly. Bugs leading to unintended privacy leakage, in some cases completely breaking DP guarantees, have been found in many implementations [73, 123, 57, 70, 9, 7, 48]. For instance, Lyu et al. [73] and Zhang et al. [123] focus on the differentially private Sparse Vector Technique (SVT) [41] and show, through manual analysis, that most of its implementations [100, 42, 68, 106, 24, 73] are not differentially private at all.
@@
However, this approach faces some fundamental limitations. First, type systems implicitly trust that some “basic” DP algorithms are implemented correctly by the underlying programming language (e.g., Gaussian mechanism, Laplace mechanism). As pointed out in [70], this may not necessarily be the case if, e.g., the noise addition mechanisms suffer from floating-point vulnerabilities [83]. Second, these type systems require mechanisms to be implemented in a very specific language, thus, they are substantially limited to simple algorithms like SVT or Report Noisy Max [42]. Specifically, they struggle to capture the complexities involved in algorithms like RAPPOR [44] or DP-SGD [1],
DP Auditing’s Role. Rather than statically checking program specifications, DP auditing allows an auditor to run experiments and statistically refute the theoretical privacy guarantees provided by DP. In other words, if the auditor fails, programmers and clients can be assured that the implementation satisfies the DP guarantees at some (possibly pre-defined) level of confidence—e.g., for more critical applications, the auditor can perform the test at 99.999% confidence [109] or 95% for less critical applications [55]. Overall, DP audits can offer more flexibility than static methods, accommodating more complex algorithms such as private machine learning pipelines [55]. They are also easier to implement as they do not involve specific programming languages or typing rules.
Empirical Privacy Estimation. DP auditing can also provide precise estimates of the privacy leakage in different settings, e.g., vis-à-vis different real-world adversaries and/or attacks.
@@ 3.2 DP Auditing Algorithms
The above definition also encodes a notion of “correctness” for DP auditing algorithms, which informally states that the output of the algorithm agrees with the actual guarantees satisfied by the mechanism (i.e., if satisfies -DP, the algorithm confirms this to be true as well). However, this definition admits vacuous algorithms (e.g., an algorithm that always outputs ‘1’); thus, in Section 3.3, we define additional desirable properties of DP auditing algorithms.
@@
First, the Adversary chooses an initial dataset in the data domain and distributions from which special data points used to audit the mechanism, i.e., canaries, are sampled from. They can also provide “auxiliary” information to the mechanism if necessary.
Next, the Challenger runs the mechanism times, each time sampling fresh canaries from the canary distributions to form the input dataset. The Challenger then sends the outputs back to the Adversary, who post-processes the outputs of the mechanism, typically using a privacy attack . This is because the outputs of the mechanism may be in an arbitrary domain , which can be difficult to design a decision function around (e.g., to decide if the mechanism satisfies the claimed guarantees). Thus, the post-processing yields simple observations that can be easy to analyze.
Finally, the outcome of the audit can be computed from the observations using an evaluation function , which outputs either a bit for the verification of a theoretical DP guarantee or for empirical privacy estimation. This is typically done by relying on the hypothesis testing interpretation of DP discussed above. In practice, the outcome of the audit is typically computed with a desired level of confidence , especially when refuting/confirming the claimed DP guarantees of the mechanism.
@@ 3.3 Desiderata: The Gold Standard of DP Auditing
Next, we define three key properties we believe audits should satisfy – efficiency, end-to-end-ness, tightness. We believe that achieving these three properties simultaneously would be optimal for the widespread and rigorous adoption of DP auditing, especially in production systems, although this remains an open research problem. Even when they cannot all be achieved simultaneously, we believe that analyzing DP audits along these three dimensions is crucial in helping data and privacy practitioners choose the optimal auditing techniques for their specific contexts.
@@
A DP auditing algorithm is efficient if it requires the mechanism to be run only a few times222Although “few” can be context-dependent, in this work, we consider “few” to mean once or twice..
End-to-End. DP audits should also apply to end-to-end privacy guarantees of the implementations; ideally, these should be run by simply providing inputs to the mechanism and observing its final output. This is often referred to as “black-box” auditing [69] and reflects threat models in practical settings as opposed to worst-case threat models assumed by the privacy analysis of certain mechanisms. For instance, the privacy analysis of the DP-SGD [1] algorithm assumes that all intermediate models are released, even though in practice only the final model may be accessible to the adversary. Since “black box” may have different meanings in different areas (e.g., query-access to machine learning models [104] vs. access to fixed generated synthetic data records [53], etc.), we use the notion of “end-to-end” auditing to avoid confusion.
End-to-end audits are also broadly applicable to many mechanisms as they do not require access to how the mechanism is implemented internally. For example, although different DP mechanisms can be used to train private ML models (e.g., DP-SGD [1], PATE [94], DP-ZO [108]), relying on one single auditing technique that takes in input the training dataset and observe the trained model’s outputs would make it significantly easier to integrate auditing into different training pipelines. Once again, this would also more closely match the capabilities of realistic adversaries.
@@
Tightness. Informally, a tight audit means that the empirical audit is consistent with the theoretical privacy analysis. More precisely, the empirical estimates obtained through auditing should not be far from the theoretical guarantees (e.g., and ) – see Definition 3.5.
When verifying DP guarantees, tight DP audits ensure that implementations are verified only if their actual privacy leakage is less than or equal to the theoretical DP bounds. Consequently, tight audits are required in this context to ensure accuracy and provide strong confidence in the process. By contrast, a loose audit that successfully verifies implementations in which the actual privacy leakage is, say, twice the theoretical upper bound would inherently not be very useful for identifying bugs.
In theory, the definition below can be proved asymptotically or demonstrated empirically. All papers we consider generally report empirical estimates of -DP, along with a confidence interval. Therefore, we consider an audit “tight” if the theoretical guarantee lies within the confidence interval of the estimate with a “sufficiently small” confidence interval.
@@
4 The Current Body of Work on
DP Auditing
We now present our methodology to compile a comprehensive and representative snapshot of DP auditing research.
@@ 4 The Current Body of Work on
Final List. After manual review, we identified 45 relevant DP auditing papers, which form the basis for our systematization effort. (Note that prior SoKs in adjacent fields typically systematize 25 to 30 papers [37, 91, 81].) To ease presentation, in Table 1, we report 21 key papers (out of the 45) that arguably represent the state of the art, while reporting earlier/seminal work in Appendix B. Specifically, Table 1 lists papers that are the first to audit a particular DP guarantee (e.g. [34, 87]), mechanism (e.g. [55, 76]), or achieve the current “best” auditing results (e.g., [87, 116]) and are typically published at top-tier venues. Along with the relevant selected papers, we also report key details of the auditing methods introduced in these papers, which we extract through our systematization process as discussed later in Section 5.
At first glance, Table 1 reveals that state-of-the-art DP auditing spans a wide range of mechanisms, from “simple” ones (e.g., Laplace, Noisy Histogram, etc.) to more “complex” machine learning pipelines, vis-à-vis various DP guarantees and adjacency notions and employing different sub-sampling techniques. Consequently, to support this wide range of mechanisms and techniques, auditing algorithms use different privacy attacks and evaluation functions, which we analyze in Sections 5.2 and 5.3, respectively. Through our systematization, we find that DP auditing techniques have met the three desiderata discussed in Section 3.3 only for simple mechanisms (e.g., Gaussian), whereas achieving them for complex mechanisms remains elusive.
Although our selection was done in January 2025, we are confident it is meaningfully representative of the state of the art in DP auditing as of the time of submission; i.e., even if one were to add or remove a small number of papers to our list, our main takeaways would not substantially change.
Out-of-scope Research. As our systematization focuses on DP auditing, we do not include work estimating privacy leakage using privacy metrics other than DP guarantees, e.g., using attack success rate, TPR@low FPR, and/or adversarial advantage [97, 56, 20, 3, 121, 114]. We also excluded research focusing on using static methods, such as type checkers, to verify the DP guarantees of an implementation without executing the program. While these methods are useful in various ways, they are arguably limited in the range of mechanisms they can audit, which makes them ineffective for verifying the DP guarantees of complex modern DP mechanisms, e.g., in the context of deep learning, as discussed in Section 3.1.
DP-SGD. Although in recent years, much of the focus on DP auditing has been on DP-SGD [1], our systematization takes a broader view of DP auditing as even for “simple” mechanisms, tight, end-to-end, and efficient have only been shown to be possible recently [75, 116]. Furthermore, many of the novel audit methodologies that have been introduced for DP-SGD are modular in principle and can be applied to any mechanism in general [105, 75, 116]. Therefore, in this work, we derive findings and recommendations that are broadly applicable to any DP mechanism in general.